OpenStudy (anonymous):

Solve using the quadratic formula. Approximate answers to the nearest tenth. 3x^2+2x-5=0

OpenStudy (anonymous):

given the form \[ax^2+bx+c=0\] the quadratic formula is \[x= \frac{ -b\pm \sqrt{b^2-4ac}}{2a}\] i suggest you memorize this formula

OpenStudy (anonymous):

I do have it memorized

OpenStudy (anonymous):

then what is your issue?

OpenStudy (anonymous):

But just because I have it memorized doesn't mean I got the question right. So im trying to see what I did wrong. That's my issue.

OpenStudy (anonymous):

well a= ? b= ? c= ?

OpenStudy (anonymous):

Are you here to help me && show me? Or just make me feel stupid?

OpenStudy (anonymous):

no, im trying to see if you made a mistake when calculating

OpenStudy (anonymous):

or when determining a, b, and c

OpenStudy (jim766):

\[\frac{ -2\pm \sqrt{(2^2)- 4(3)(-5)} }{ 6 }\] is this how you set it up?

OpenStudy (anonymous):

yeah